Lena plans to invest 8,700 dollars in 2 year(s) and 8,000 dollars in 1 year(s). She expects to earn 15.24 percent, compounded quarterly. How much money does Lena expect to have in 10 years?
Answer #1
Step-1:Calculation of future value of 8,700
Future value = P * (1+i)^n Where,
= $   8,700.00 * (1+0.0381)^32 P = $   8,700.00
= $   8,700.00 * 3.30867503 i = 15.24%/4 = 0.0381
= $ 28,785.47 n = 8*4 = 32
Step-2:Calculation of future value of 8,000
Future value = P * (1+i)^n Where,
= $   8,000.00 * (1+0.0381)^36 P = $   8,000.00
= $   8,000.00 * 3.84247347 i = 15.24%/4 = 0.0381
= $ 30,739.79 n = 9*4 = 36
Step-2:Calculation of value of money in 10 years
Value of money in 10 years = $ 28,785.47 + $ 30,739.79
= $ 59,525.26